OpenStudy (anonymous):

you just bought 4 pairs of jeans for $32 write a ratio and sentence comparing the number of jeans bought to the cost

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

since 4 jeans cost $32 the price per jean is \(\frac{32}{4}=8\) i.e. $8 per jean (pretty cheap) the other way \(\frac{4}{32}=\frac{1}{8}\) means your dollar bought one eights of a pair of jeans